OpenStudy (anonymous):

-6p+7=3(2p-3)-4(-10+4p) what is p a)p=6 b)p=5 c)p=7 d)p=12

OpenStudy (texaschic101):

do you know how to distribute through the parenthesis ?

OpenStudy (anonymous):



OpenStudy (texaschic101):

check your math again

OpenStudy (texaschic101):

@herbfu ...check your math again

OpenStudy (anonymous):

sorry that it took so long to reply i had to put up my lawn mower

OpenStudy (anonymous):

The first thing you need to do is simplify the expressions on both sides. Distribute if you need to after. After this, you need to use the addition and subtraction properties of equality to get variables on one side of the equation, and constants on the other. Repeat if necessary. After this, use the division property of equality to get the final answer to your equation.

OpenStudy (anonymous):

still dont understand

OpenStudy (texaschic101):

Let me show you... -6p + 7 = 3(2p - 3) -4(-10 + 4p) -6p + 7 = 6p - 9 + 40 - 16p (combine like terms) -6p + 7 = - 10p + 31 (add 10p to both sides) -6p + 10p + 7 = 31 (subtract 7 from both sides) -6p + 10p = 31 - 7 (combine like terms) 4p = 24 (divide both sides by 4) p = 6 check.. -6(6) + 7 = 3(2(6) - 3) -4(-10 + 4(6) -36 + 7 = 3(12 - 3) - 4(-10 + 24) -29 = 3(9) - 4(14) -29 = 27 - 56 -29 = -29 (correct) p = 6

OpenStudy (anonymous):

-6p+7=3(2p-3)-4(-10+4p) Distribute the 3 across the first parentheses, and the 4 across the second, giving you -6p+7=6p-9+40-16p Combine your like terms. -6p+7=-10p+31 Now, solve finish solving the equation using the addition, subtraction, and division properties of equality.